Passing through the point (2,8) and parallel to the line equation 3x+2y=13
Determine the equation of the line

Answer by ewatrrr(24785)

 
Hi,


Passing through the point (2,8) and parallel to the line equation 3x+2y=13
parallel to the line equation
3x+2y=13 OR y = (-3/2)x + 13/2 m = (-3/2)
Parallel lines have identical slopes
New Line
y = (-3/2)x + b |Using ordered pair (2,8) to solve for b
8 = (-3/2)*2 + b
11 = b
y = (-3/2)x + 11 OR 3x+2y = 22